Using Extended Euclid

a. Use Euclid’s algorithm to compute gcd(1175, 423)

b. Use the extended Euclid algorithm to find integers x and y such that gcd(1175, 423) = 1175x + 423y. What is x and y?
Answer #1
a)
=> GCD(1175, 423)
=> GCD(423, 1175%423)
=> GCD(423, 329)
=> GCD(329, 423%329)
=> GCD(329, 94)
=> GCD(94, 329%94)
=> GCD(94, 47)
=> GCD(47, 94%47)
=> GCD(47, 0)
GCD(1175, 423) is 47

b)
gcd(1175, 423) = 47
we can write 1175 as 47*25
we can write 423 as 47*9

gcd(1175, 423) = 1175/(2*25) + 423/(2*9)
x = 1/50
y = 1/18